OpenStudy (anonymous):

(x^2+5)(square root of x-3)

OpenStudy (anonymous):

what are you supposed to do with it?

OpenStudy (anonymous):

\[(x^2+5)/\sqrt{x^3}\] \[x^4+25/x^3\] i just tried am not certain

OpenStudy (anonymous):

multiply them

OpenStudy (anonymous):

\[(x^2+5)(\sqrt{x}-3)\]like this?

OpenStudy (anonymous):

yeah but the square root is over the 3 too

OpenStudy (anonymous):

\[(x^2+5)(\sqrt{x-3})\]like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so not much you can do. you can write \[x^2\sqrt{x-3}+5\sqrt{x-3}\] if you like. just the distributive law. cannot do anything else really

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

isn't sqrt(x-3) same thing as (x-3)^1/2 ?

OpenStudy (anonymous):

yeah